# Hierarchical Clustering Algorithms on Poisson and Cox Point Processes

### v2

@@ 5.1 Proof of $\mathcal{F/F}$ property of the pre-limits point-shift graphs of $\mathrm{CHN}^{2}$
5.2 Proof of Theorem 3.2
To prove Theorem 3.2, an approach similar to that of the proof of Corollary 2.8 is proposed. We recall that the proof of Corollary 2.8 relies on the fact that the intensity of the exit points at level $k$ of the $\mathrm{CHN}^{2}$ clustering algorithm, tends to zero as $k\to\infty$. For the weak convergence of the graphs $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$ constructed at level $k$ in the $\mathrm{SHN}^{2}$ clustering algorithm, a similar argument is required. Specifically, it must be shown that the intensity of proximity points of level $k$, denoted by $\lambda_{k}$, also tends to zero as $k\to\infty$. To establish this, this section defines a sequence of vertex shifts, $f^{k}_{s}$, for $k\geq 0$, on the random graphs $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$. The definition of these vertex shifts is based on the notion of mutual nearest neighbor pairs, which are well defined within each finite cluster. Head points and exit points are defined as in Section 2 for the $\mathrm{CHN}^{2}$ algorithm. Finally, using the intensity of exit points, it is shown that the intensity of proximity points at each level tends to zero as $k\to\infty$.
It is important to note that the vertex shift of order $k$ does not coincide with the edges of the graph at different levels of the algorithm. In particular, the vertex shift graph on $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$ is only used for assessing $\lambda_{k}$, and it should not be confused with the undirected graph of interest here.
Order 0
[…]
Each connected component of the $f^{0}_{s}$-graph on the PPP contains a cycle of length two. These cycles are again referred to as 0-cycles. The vertices in the $f^{0}_{s}$-graph that belong to the 00-cycles are still called cluster heads of order zero. The union of cluster heads of order zero forms a thinning of $\Phi^{0}$, denoted by $\Phi^{0}_{\mathcal{S},c}$. Let $\Phi^{s,0}_{a}=\Phi^{0}\setminus\Phi^{s,0}_{c}$, which denotes points not in 00-cycles.
Order 1
To define the second point-shift, mutual nearest neighbors within the clusters of order zero are used. The definition is similar to the $\mathrm{CHN}^{2}$ clustering algorithm on PPP, but instead of using the distance between mutual nearest neighbor pairs within clusters, we use the distance between clusters, specifically with respect to the single linkage distance. For this, we need the following definition:
Definition 5.9.
Let $\{C^{0}_{i}\}_{i}$ denote the collection of 00-cycles, and let $\{S^{0}_{i}\}_{i}$ be the collection of pairs of points in these cycles. Specifically, $S^{0}_{i}$ is the pair of head points belonging to the cycle $C^{0}_{i}$, and $\mathcal{C}^{0}_{i}$ is the cluster in $[\mathcal{G}^{1}_{\mathcal{S}},0]$ that $C^{0}_{i}$ belongs to. The sets $\{S^{0}_{i}\}_{i}$ form a translation-invariant partition of the support of $\Phi^{s,0}_{c}$.
Now, consider the nearest neighbor map, denoted by NN1, on $\{S^{0}_{i}\}_{i}$. This map associates each pair $S^{0}_{i}$ with $S^{0}_{j}$, where $S^{0}_{j}$ belongs to $\mathcal{C}^{0}_{j}$, and $\mathcal{C}^{0}_{j}$ is the nearest neighbor of $\mathcal{C}^{0}_{i}$ based on the single linkage distance.
[…]
Define $f^{1}_{s}$ to be the vertex shift on the vertices of $[G^{1}_{S},0]$, where $f^{1}_{s}$ coincides with $f^{0}_{s}$ everywhere except at the exit points of order 0. For each such point, say $x\in S^{0}$, where $S^{0}$ is the pair in which $x$ is the exit point, define $f^{1}_{s}(x)=\mathrm{nn1}(x)$.
[…]
Definition 5.10.
[…]
If the edges of the $1$-cycles in the $f^{1}_{s}$-graph are removed, the remaining structure consists of directed trees, which we call cluster subtrees of order 1. Each vertex $x$ in a cluster subtree of order 1 is either a cluster head of order 1 (i.e., $x\in\Phi^{s,1}_{c}$) or an element of $\Phi^{s,1}_{a}$, with a directed path leading from $x$ to a cluster head of order 1. This structure is analogous to the point-shift graph of order 1 in the $\mathrm{CHN}^{2}$ clustering algorithm on Poisson point processes.
Order 2
To define the next vertex shift, $f^{1}_{s}$, on the graph $[\mathcal{G}^{2}_{\mathcal{S}},0]$, we need to introduce some concepts similar to those in Definition 5.9.
Definition 5.11.
Let $\{C^{1}_{i}\}_{i}$ denote the collection of $1$-cycles, and let $\{S^{1}_{i}\}_{i}$ represent the collection of point pairs in these cycles. Specifically, $S^{1}_{i}$ is the pair of head points in the cycle $C^{1}_{i}$, and $\mathcal{C}^{1}_{i}$ is the connected component within $[\mathcal{G}^{2}_{\mathcal{S}},0]$ to which $C^{1}_{i}$ belongs. The sets $\{S^{1}_{i}\}_{i}$ form a translation-invariant partition of the support of $\Phi^{s,1}_{c}$.
Now, consider the nearest neighbor map, denoted by NN2, on $\{S^{1}_{i}\}_{i}$. This map associates each pair $S^{1}_{i}$ with $S^{1}_{j}$, where $S^{1}_{j}$ belongs to $\mathcal{C}^{1}_{j}$, and $\mathcal{C}^{1}_{j}$ is the nearest neighbor of $\mathcal{C}^{1}_{i}$ based on the single linkage distance.
[…]
Define $f^{1}_{s}$ as the vertex shift on $[\mathcal{G}^{2}_{\mathcal{S}},0]$. Note that $[\mathcal{G}^{2}_{\mathcal{S}},0]$ coincides with $[\mathcal{G}^{1}_{\mathcal{S}},0]$ except for the proximity points of each connected component of $[\mathcal{G}^{1}_{\mathcal{S}},0]$. Define $f^{1}_{s}$ coincides with $f^{1}_{s}$ everywhere except for the exit points of order 1. For each such point, say $x\in S^{2}$, where $S^{2}$ is the pair in which $x$ is the exit point, define $f^{2}(x)=\mathrm{nn2}(x)$. Again, $f^{2}$ defines a vertex shift on $[\mathcal{G}^{2}_{\mathcal{S}},0]$. The connected components of the $f^{2}$-graph are identical to those of $[\mathcal{G}^{2}_{\mathcal{S}},0]$, though the edges do not coincide everywhere. As mentioned earlier, without loss of generality, we assume that the connected components of the $f^{1}_{s}$-graph may be finite or infinite with positive probability. The finite connected components contain a unique cycle of length 2 almost surely, as ties do not occur almost surely. The points within these 2-cycles are exit points of order 1 and form a sub-Poisson point process (sub-PP) of $\Phi^{s,1}_{c}$, called the cluster heads of order 2 and denoted by $\Phi_{c}^{S,2}$. Let $\Phi_{a}^{S,2}:=\Phi^{0}\setminus\Phi_{c}^{S,2}$, representing the vertices in the acyclic part of the $f^{2}$-graph.
Order k
Similarly, we can define, by induction, analogous notions for each level $k\geq 2$:
As illustrated in the definitions of point-shifts $f^{1}_{s}$ and $f^{1}_{s}$, the point-shift $f^{k}_{s}$ at level $k$ can be constructed using the previous level’s point-shift, $f^{k-1}_{s}$. By induction, for each level $k$, we define the following notions:
$k$-cycles
$\{C^{k}_{i}\}_{i}$, consisting of connected components of $f^{k}_{s}$ graphs that contain cycles of length 2;
Pairs within $k$-cycles
$\{S^{k}_{i}\}_{i}$, where each pair in $\{S^{k}_{i}\}_{i}$ represents the head points of the cycle $C^{k}_{i}$ in the $f^{k}_{s}$-graph;
Nearest Neighbor Map, NNk
: for each pair $S^{k}_{i}$, $\mathrm{NNk}(S^{k}_{i})$ denotes the nearest neighboring pair $S^{k}_{j}$ in the single linkage sense, where $S^{k}_{j}$ belongs to the nearest neighbor cluster $\mathcal{C}^{k}_{j}$ of $\mathcal{C}^{k}_{i}$;
Nearest Neighbor Point Map, $\mathbf{nnk}$
: for each pair $S^{k}_{i}$, $\mathrm{nnk}(x_{i})=x_{j}$ identifies the nearest neighboring point $x_{j}\in S^{k}_{j}$ for a point $x_{i}\in S^{k}_{i}$, achieving the minimum distance between the two pairs, when $\mathrm{NNk}(S^{k}_{i})=S^{k}_{j}$.
Points within these $k$-cycles which are exit points of order $k-1$ form a sub point process (sub-PP) of $\Phi^{s,k}_{c}$, referred to as cluster heads of order $k$ and denoted by $\Phi_{c}^{s,k}$. Let $\Phi_{a}^{s,k}:=\Phi^{0}\setminus\Phi_{c}^{s,k}$ represent the vertices in the acyclic part of the $f^{k}_{s}$-graph.
Remark 5.12.
Let $\Phi^{s,k}_{e}$ be the point process of exit points of order $k$, obtained by applying the vertex shift $f^{k}_{s}$. Suppose the intensity of $\Phi^{s,0}_{e}$ is $\rho$. By the same argument given in Remark 2.6, one can show that the intensity of $\Phi^{s,k}_{e}$ is less than $\rho/2^{k}+1$. Consequently, the weak limit graph of the vertex-shift graphs ($f^{k}_{s}$-graphs) exists and is denoted by the $f^{\infty}_{s}$ graph.
Proof of Theorem 3.2.
Consider the $\mathrm{SHN}^{2}$ (or $\mathrm{H}^{2}\mathrm{N}^{2}$) algorithm on $\Phi^{0}$. By the definition of $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$ and of $f^{k}_{s}$-graphs, each connected component of $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$ contains exactly the same vertices as the corresponding connected component of the $f^{k}_{s}$-graph, and vice versa. That is, each connected component $C_{k}$ in $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$ can be mapped in a bijective way to one connected component $C^{\prime}_{k}$ in the $f^{k}_{s}$-graph.
Similarly, each proximity point in $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$ can be mapped bijectively to one exit point in the $f^{k}_{s}$-graph, and vice versa. Hence, the intensity $\lambda_{k}$ of the proximity points in $[\mathcal{G}^{k}_{\mathcal{{S}}},0]$ equals the intensity of the order-$k$ exit points in the $f^{k}_{s}$-graph. By Remark 5.12, one concludes that $\lambda_{k}<\rho/2^{k}$.
Since $\lambda_{k}\to 0$ as $k\to\infty$, the same argument as the one used in the proof of Corollary 2.8 applies here to show that the weak limit of the graphs $[\mathcal{G}^{n}_{\mathcal{{S}}},0]$ on $\Phi^{0}$ exists as $n\to\infty$. ∎
6 Connection with Other Spanning Forests on Poisson Point Process
@@
In this paper a hierarchical clustering model is introduced using the $\mathrm{CHN}^{2}$ clustering algorithm, which creates a sequence of random graphs in a Poisson point process (PPP). Under Palm probability, its weak limit graph, $\mathrm{CHN}^{2}$ Spanning Forest, exists. The $\mathrm{CHN}^{2}$ Spanning Forest is a forest on PPP, where each connected component of it is one-ended. This section aims to compare the similarities and differences between this model and other known structures in the literature constructed on PPP.
@@
The first construction related to $\mathrm{CHN}^{2}$ Spanning Forest is the factor graphs on PPP. Ferrari, Landim, and Thorisson [14] provide a deterministic algorithm to construct a graph with no loops (either a tree or a forest), where the vertices are the points of a d-dimensional stationary Poisson process. For dimensions 2 and 3, their algorithm results in a one-ended tree. Holroyed and Peres [17] shown that in any dimension $d$, a factor graph of the Poisson point process $\Phi^{0}$ can be constructed as a one-ended tree. A factor graph on a point process is a random graph $G$ whose vertex set is the support of the point process, and $G$ is a deterministic function invariant under isomorphisms on the PPP. In this sense, the $\mathrm{CHN}^{2}$ EFF graph $[G,0]$ defined on the PPP can also be seen as a factor graph, forming a one-ended forest.
In their proof, Holroyed and Peres construct different levels $n\in\mathbb{N}$ of coarser partitions on the PPP, resulting in infinitely many finite connected components at each level. Each component at level $n$ belongs to exactly one component at level $n+1$, and in the limit, there remains only one connected component. Using this structure, they define a one-ended tree on the PPP. While their construction involves a hierarchical structure similar to ours, it is analytical and does not provide an explicit algorithm for constructing the different levels of connected components.
@@
As an extension, Timar [25] showed that the existence of a tree factor graph on the PPP can be extended to any stationary non-equidistant point process.
Another well-known model is the Minimal Spanning Forest (MSF). Aldous and Steele [23] introduced the MSF as a natural extension of the Minimal Spanning Tree (MST) to infinite graphs and point processes. They conjectured that for the PPP, their MSF consists of a single one-ended tree. However, it remains generally unknown how many components this algorithm produces and how many ends each component has (some results for dimension 2 are found in Alexander’s work [3]). Notably, the first level of the $\mathrm{CHN}^{2}$ clustering algorithm, i.e., $[G^{0},0]$, forms a subgraph of the MSF on the PPP. However, from the second level onward, due to the algorithm’s greedy nature, significant differences emerge.
Finally, a model related to random thinning, which naturally leads to a classification algorithm, is another relevant structure. Consider a stationary point process $\Phi$. Let $\pi=\{p_{k}\}$ be a probability distribution on the positive integers. Independently mark the points of $\Phi$ according to $\pi$. Then, perform a random thinning of $\Phi$, retaining only points of $\Phi^{0}$ marked by $k$. Declare the retained points $\Phi^{k}_{c}$ to be the cluster heads of level $k$, and assign as descendants the points of $\Phi^{k-1}_{c}$ that lie within the Voronoi cell of their nearest cluster head in $\Phi^{k}_{c}$. This algorithm, studied in [27], requires additional randomness, while clustering in $\mathrm{CHN}^{2}$ does not.
7 Open Problems
Several open questions and conjectures remain regarding the $\mathrm{CHN}^{2}$ point-shift and the $\mathrm{CHN}^{2}$ EFF. One such question revolves around whether the $\mathrm{CHN}^{2}$ EFF is a tree when constructed on the PPP. The conjecture is that, in all dimensions, this point-shift graph is a tree on a PPP.
Consider a point-shift $f$ on a point process (PP) and the action of the semigroup of translations by $-f$ on probability distributions of counting measures that have a point at the origin. The $f$-probability (point map probability) of the point process is then defined as the limit of the orbit of this semigroup action on the Palm distribution of the PP when it exists. Another question related to $\mathrm{CHN}^{2}$ point-shift is whether the point map probability of this point-shift exists on the PPP, and how to compute it if it exists. Answering these questions and validating the conjecture would greatly contribute to our understanding of the algorithm’s behavior.
In the context of the $\mathrm{SHN}^{2}$ and $\mathrm{H}^{2}\mathrm{N}^{2}$ clustering algorithms on Poisson point processes (PPP), a number of questions remain open as well:
  • In Corollary 3.4, it is shown that on a PPP the $\mathrm{SHN}^{2}$ Spanning Forest is a subgraph of the Wired MSF. While these two graphs coincide for any finite point set, an open question is whether they are also identical on a PPP.
  • Are all clusters at different levels of the algorithm almost surely finite in $\mathrm{SHN}^{2}$ and $\mathrm{H}^{2}\mathrm{N}^{2}$ clustering algorithms?
  • How many ends does the $\mathrm{SHN}^{2}$ Spanning Forest have in theses cases?
  • How many connected components does the $\mathrm{SHN}^{2}$ Spanning Forest have in these cases? In particular, is it a single tree or a forest?

### v3

@@ 4 Proof of $\mathcal{F/F}$ property of the pre-limits graphs of $\mathrm{CHN}^{2}$
5 Related Work and Open Problems
Related Work
@@ 5 Related Work and Open Problems
This work introduces a hierarchical clustering scheme on a Poisson point process (PPP) via the $\mathrm{CHN}^{2}$ algorithm, generating a sequence of random graphs whose local weak limit under Palm probability defines the $\mathrm{CHN}^{2}$ spanning forest (EFF). Each connected component of the $\mathrm{CHN}^{2}$ EFF on the PPP is one-ended. We briefly situate this construction among related spanning structures.
Factor graphs on PPP.
@@
Ferrari, Landim and Thorisson [Hermann] gave a deterministic (factor) construction on a $d$-dimensional stationary PPP producing a loopless graph, a tree in dimensions $d=2,3$, with one end. Holroyd and Peres [Matching] further showed that in any dimension a factor of the PPP can be a one-ended tree. A factor here is a graph whose vertex set is the support of the point process and whose edge set is a deterministic, isomorphism-invariant function of the configuration. In this sense, the $\mathrm{CHN}^{2}$ EFF on the PPP is also a factor graph (constructed level by level), and it is a one-ended forest.
The Holroyd–Peres proof proceeds by building a nested, hierarchical sequence of partitions of the PPP with finitely many finite components at each level, such that every level-$n$ component is contained in exactly one level-(n+1)(n\!+\!1) component, yielding a single component in the limit; this structure then induces a one-ended tree [Matching]. While their hierarchy is conceptually close to ours, their argument is analytical and does not prescribe an explicit algorithm for forming the components at each level.
@@
Timar [timar2009] extended existence of tree factors beyond PPP to general stationary non-equidistant point processes.
Minimal Spanning Forests.
Aldous and Steele [steele_aldous1992] introduced the minimal spanning forest (MSF) as an infinite-volume analogue of the minimal spanning tree for random points, and conjectured that the MSF of the PPP is almost surely a single one-ended tree. The number of components and their ends remain subtle in general (sharp results are known in special settings; see, e.g., [alexander1995]). The level-0 graph $G^{0},0]$ produced by $\mathrm{CHN}^{2}$ is a subgraph of the MSF on the PPP; beyond level 0, the greedy merging rule of $\mathrm{CHN}^{2}$ leads to structural differences from the MSF.
Random thinning and hierarchical heads.
A thinning-based hierarchical classifier related to clustering is studied in [zuyev_tchoumatchenko2001]: points are independently marked by a distribution $\pi=\{p_{k}\}$, then thinned to retain level-$k$ heads, and lower-level points are attached within Voronoi cells of their nearest higher-level head. This scheme uses additional external randomness, whereas $\mathrm{CHN}^{2}$ is an intrinsic (deterministic) factor of the underlying point process.
Open Problems
We record several questions motivated by the $\mathrm{CHN}^{2}$ point-shift and its EFF.
  1. 1.
    Tree conjecture on PPP. Is the $\mathrm{CHN}^{2}$ EFF a tree (i.e., acyclic and connected componentwise) on the PPP in every dimension? Current results establish one-endedness of components; ruling out cycles at all scales remains open.
  2. 2.
    Point-map probability for $\mathrm{CHN}^{2}$. For a point-shift $f$ acting on configurations, consider the semigroup of translations induced by f-f on probability measures supported on configurations with a point at the origin. The $f$-probability (point-map probability) is the limit of the orbit of this action applied to the Palm distribution, when it exists (see, e.g., [Pointshift]). Does the point-map probability exist for the $\mathrm{CHN}^{2}$ point-shift on the PPP? If so, can it be computed explicitly?
  3. 3.
    Alternative inter-cluster distances. The current $\mathrm{CHN}^{2}$ hierarchy uses the clustroid distance in its merging rule. Analyze the variants obtained by replacing this distance with, for example, Hausdorff distance or single-linkage distance between clusters. For these variants, determine:
    • finiteness of connected components at each hierarchical level;
    • existence and identification of the local weak limit under Palm probability;
    • foil classification (e.g., $\mathcal{F/F}$ at finite levels and $\mathcal{I/I}$ in the limit) and the number of ends of components.
  4. 4.
    Beyond PPP. On stationary Cox or, more generally, stationary non-equidistant point processes (cf. [timar2009]), to what extent do the PPP results persist?